OpenStudy (anonymous):

Find the volume of the region between the graph of f(x,y)=9−x^2−y^2 and the xyplane

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

thank god

OpenStudy (anonymous):

i got 2 questions left and i'm done with calc 3 forever

OpenStudy (anonymous):

so xy plane is z=0 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we have to think about the region of integration

OpenStudy (anonymous):

9−x^2

OpenStudy (anonymous):

it's much easier when i can just pop it into wolfram lol

OpenStudy (anonymous):

\[\int _{-3}^3\int _{-\sqrt{9-x^2}}^{\sqrt{9-x^2}}\left(9-x^2-y^2\right)dydx\]

OpenStudy (anonymous):

holy crap you can actually do that?

OpenStudy (anonymous):

pop it into wolfram lol

OpenStudy (anonymous):

i gotta learn wolfram better

OpenStudy (anonymous):

no...just my last 2 homework questions

OpenStudy (anonymous):

test was this past monday...just some lingering homework before the final

OpenStudy (anonymous):

u got time for another if i make another post?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

I am sorry to see how many time you had to post this